OpenStudy (anonymous):

Find the derivative of f(x) = 8x^2 + 11x at x = 7.

OpenStudy (anonymous):

how do you do this?

OpenStudy (anonymous):

find f'(x) which is = 16x +11 then sub x=7 in f'(x) so answer is 16(7)+11

OpenStudy (anonymous):

using the power rule \[f'(x)=2(8)x^1+1(11)x^0 =16x+11\]

OpenStudy (anonymous):

first you need to find the derivative of f(x) = 8x^2 + 11x, to do this \[f \prime(x) = 2*(8x ^{(2-1)}) + 1*11x^{(1-1)}\] which is \[f \prime(x) = 16x + 11\] then let x = 7 gives us \[f \prime(7) = 123\]